OpenStudy (anonymous):

A point is moving along the graph of a given function such that dx/dt is 2 centimeters per second. Find dy/dt for the given values of x. y = 9x2 + 5 (a) x = -1 dy/dt =____ cm/sec (b) x = 0 dy/dt =____cm/sec (c) x = 3 dy/dt =_____ cm/sec

OpenStudy (anonymous):

\[y'=18xx'\] and you are told that \[x'=0\] so you get \[y'=36x\] just plug in the numbers

OpenStudy (anonymous):

sorry i mean you are told that \[x'=2\]

OpenStudy (anonymous):

this one is mile easier than the first ones. i wonder why they get easier.

OpenStudy (anonymous):

I have is the question I am trying to take my ASVAB test and this question I do not get at all 4! that is all the information they give

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@navyman are you asking about \[4!\] did you get lost in this post? am i hallucinating?

OpenStudy (anonymous):

so satletite what would the numbers be like hwat am i plugging in

OpenStudy (anonymous):

@lostdove all you have to so is write \[\frac{dy}{dt}=36x\] and so first one is \[36\times -1=-36\] second one is \[36\times 0=0\] third one is \[36\times 3=108\] easy an done

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

yw